# Any-Step Density Ratio Estimation via Interval‑Annealed Secant Alignment

### v1

@@ 2 Related Works
Seminal works in DRE follow two main avenues: estimating each density separately (e.g., via kernel density estimation (Huang et al. 2006)) or modeling the ratio directly by minimizing a statistical divergence, as in KLIEP (Sugiyama et al. 2008) and uLSIF (Kanamori, Hido, and Sugiyama 2009). While the latter avoids explicit density modeling, these methods struggle when the two distributions have little overlap or disjoint supports–a scenario known as the density-chasm problem (Gutmann and Hyvärinen 2012; Liu et al. 2017). To bridge this gap, TRE (Rhodes, Xu, and Gutmann 2020) proposed decomposing the ratio into a sequence of intermediate steps, each modeled by a separate network. Its continuous counterpart, DRE-$\infty$ (Choi et al. 2022), introduced a time-dependent score function defined along a continuous interpolation path. This marked a conceptual advance but introduced a major drawback: inference requires costly numerical integration of the score, often involving hundreds of evaluations per sample. Later refinements, such as diffusion-bridge interpolants (Chen et al. 2025) and conditional score matching (Yu et al. 2025), improve robustness and training efficiency. However, they do not eliminate the fundamental computational bottleneck posed by numerical integration.
@@ 5.3 Mutual information Estimation
Mutual information (MI) quantifies the statistical dependence between two random variables $\mathbf{X}\sim p(\mathbf{x})$ and $\mathbf{Y}\sim q(\mathbf{y})$, and is defined as $\text{MI}(\mathbf{X},\mathbf{Y})=\mathbb{E}_{p(\mathbf{x},\mathbf{y})}\!\biggl{[}\log\frac{p(\mathbf{x},\mathbf{y})}{p(\mathbf{x})\,q(\mathbf{y})}\biggr{]}$. Since this formulation involves the expectation of a log density ratio, MI estimation naturally reduces to a DRE problem. To evaluate the robustness of our method, we introduce two challenging MI estimation benchmarks: (1) geometrically pathological distributions, where $(\mathbf{X},\mathbf{Y})$ lie on complex non-linear manifolds; and (2) high‑discrepancy Gaussian pairs, with strong correlations and anisotropic covariance structures.

### v2

@@ 2 Related Works
Seminal works in DRE follow two main avenues: estimating each density separately (e.g., via kernel density estimation (Huang et al. 2006)) or modeling the ratio directly by minimizing a statistical divergence, as in KLIEP (Sugiyama et al. 2008) and uLSIF (Kanamori, Hido, and Sugiyama 2009). Discriminative approaches such as NCE (Gutmann and Hyvärinen 2012) and TR-DRE (Liu et al. 2017) have been widely adopted, with further theoretical insights into truncation given by Song and Ermon (2019). Despite their simplicity, these methods often fail under the density-chasm problem, where $p_{0}$ and $p_{1}$ have little or no overlap, leading to unstable training and poor generalization in high-dimensional settings. To bridge this gap, divide-and-conquer strategies have been proposed. TRE (Rhodes, Xu, and Gutmann 2020) decomposes the ratio into a sequence of intermediate steps, each estimated separately, and AM-DRE (Xiao and Han 2022) adaptively refines this idea for latent-space EBM learning. Their continuous counterpart, DRE-$\infty$ (Choi et al. 2022), introduces a time-dependent score function defined along a continuous interpolation path, marking a conceptual shift from static ratios to dynamical estimation. However, these methods rely on expensive numerical integration, often requiring hundreds of evaluations per sample. Subsequent refinements explored alternative parameterizations and regularization schemes: FDRE (Choi, Liao, and Ermon 2021) learns featurized ratios to improve generalization; IMDRE (Kimura and Bondell 2025) combines importance sampling with multi-step estimation; D3RE (Chen et al. 2025), $\gamma$-DRE (Nagumo and Fujisawa 2024), and iterated regularization (Gruber et al. 2024) further enhance robustness under large-support mismatch. Srivastava et al. (2023) reformulate DRE as multinomial logistic regression to better handle high-discrepancy distributions, and Williams et al. (2025) exploit exponential-family representations to reduce variance. Despite these advances, the computational bottleneck posed by numerical integration persists, motivating approaches like ours that preserve the continuity and robustness of DRE-$\infty$ while eliminating solver overhead.
@@ 5.3 Mutual information Estimation
Mutual information (MI) quantifies the statistical dependence between two random variables $\mathbf{X}\sim p(\mathbf{x})$ and $\mathbf{Y}\sim q(\mathbf{y})$, and is defined as $\text{MI}(\mathbf{X},\mathbf{Y})=\mathbb{E}_{p(\mathbf{x},\mathbf{y})}\left[\log\frac{p(\mathbf{x},\mathbf{y})}{p(\mathbf{x})q(\mathbf{y})}\right]$. Since this formulation involves the expectation of a log density ratio, MI estimation naturally reduces to a DRE problem. To evaluate the robustness of our method, we introduce two challenging MI estimation benchmarks: (1) geometrically pathological distributions, where $(\mathbf{X},\mathbf{Y})$ lie on complex non-linear manifolds; and (2) high‑discrepancy Gaussian pairs, with strong correlations and anisotropic covariance structures.